A passenger train leaves Norfolk, Virginia, 1.2 hours after a freight train leaves. The passenger train is traveling 18 mph faster than the freight train on a parallel set of tracks. The trains are traveling IN THE SAME DIRECTION. The two trains will be at the same location 3 hours after the passenger train leaves. Find the speed of each train.
Found 2 solutions by haileytucki, stanbon:
Answer by haileytucki(390)

Rate*Time=Distance
Freight train is train 1, passenger train is train 2.
R1+T1=D
R2+T2=D
Both go to the same destination, travel the same distance
R2=18+R1
Train 2 is 18 mph faster than Train 1.
T1=4.2
T2=3
Train 2 takes 3 hours, train 1 1.2 hours longer.
R(1)*4.2=(18+R(1)*3)
4.3R(1)=54+3R(1)
1.2R1=54
R1=45Mph
R2=45+18=63 MPH

Answer by stanbon(75887)

A passenger train leaves Norfolk, Virginia, 1.2 hours after a freight train leaves.
------------------------------
The passenger train is traveling 18 mph faster than the freight train on a parallel set of tracks.
------------------------------
The trains are traveling IN THE SAME DIRECTION. The two trains will be at the same location 3 hours after the passenger train leaves.
------------------------------
Find the speed of each train.
--------------------------------------
Passenger train DATA:
time = 3 hrs ; rate = x+18 mph ; distance = 3*(x+18) = 3x+54 miles
-----------------------
Freight train DATA:
time = 4.2 hrs ; rate = x mph ; distance = 4.2x miles
--------------------------
Equation:
distance = distance
3x+54 = 4.2x
1.2x = 54
x = 45 mph (freight train rate)
---
x+18 = 63 mph (passenger train rate)
======================================
Cheers,
Stan H.